YOUSSEFben:

The function M(s) = 225 + 0.65s represents the material cost of manufacturing gardening scissors when s scissors are produced. The function L(s) = 54 + 1.15s represents the labor cost for producing s scissors. Which expression correctly represents the manufacturing cost per scissors? A. B. C. D.

AZ:

Based on the limited information provided, I assume manufacturing costs = material cost + labor cost and as such, if the manufacturing cost is C(x) then C(x) = M(x) + L(x) and so to find your answer choice, you just have to add both functions C(x) = 225 + 0.65s + 54 + 1.15s C(x) = ??